OpenStudy (anonymous):

An isosceles triangle has a base that is 5 inches shorter than either of the equal sides. If the perimeter of the triangle is 34 inches, what is the length of each side?

OpenStudy (anonymous):

equalsides = x base = x-5 x+ x + x-5 = 34 3x = 39 x = 13 x-5 = 8 three sides is 13 13 8